OpenStudy (anonymous):

PLEASE HELP!!!! 6 (9/20*2/5*3/20)

OpenStudy (anonymous):

I didn't get it write using equation

OpenStudy (anonymous):

I dont know. All it says is solve in fraction form and then gives this problem

OpenStudy (anonymous):

okay is it \[6(\frac{ 9 }{ 20}*\frac{ 2 }{ 5 }* \frac{ 3 }{ 20 })\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

lol you just have to multiply 6*9*2*3 and divide it by 20*5*20

OpenStudy (anonymous):

so would it be 81/250 or 81/500 as the answer?

OpenStudy (anonymous):

uh you are supposed to calculate fractional value

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

Okay but they give me options of 27/1000, 81/1000, 8/500, or 81/250

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Do you know know how to solve it or not?

OpenStudy (anonymous):

ah fine let me calculate for you

OpenStudy (anonymous):

81/500

OpenStudy (anonymous):

Thank you